OpenStudy (anonymous):

Simplify the expression by using a double angle formula 1-2sin^2 2pi/9

OpenStudy (jmark):

cos 2A = 1 - 2sin^2 A. Hence, 1 - sin^2 (2pi/9) = cos 2(2pi)/9 = cos (4pi/9).